# Fine-Tuning Flow Matching via Maximum Likelihood Estimation of Reconstructions

### v1

@@ 3.2 MLE Fine-tuning
The following theorem provides a concrete loss function for MLE fine-tuning.
Theorem 2.
Under Assumption 1, and when $\Sigma$ is a scalar matrix, performing maximum likelihood estimation (MLE) by maximizing the expectation $\mathbb{E}_{q(x_{0},x_{1})}\left[\log p_{1}\left(\hat{\phi}_{N}(x_{0})|x_{1}\right)\right]$ is equivalent to minimizing the following loss function
@@
The proof, along with the loss function in the more general case of a diagonal matrix $\Sigma$, is given in Appendix B. (6) provides us with a computationally feasible loss function that we can directly use to fine-tune our model. As established by Theorem 2, our method enables the direct optimization of the model error $\varepsilon_{N}(x_{0})$. This stands in contrast to previous approaches ((4) and (5) in Theorem 1), which could only provide a loose upper bound that contained non-optimizable components.
@@
MLE by (6) has the advantange of high-precision for it directly optimizing the object obtained by inference procedure. But it suffers from several critical issues. Firstly, it is particularly prone to overfitting. Since no additional constraints are imposed on the vector field or trajectory shapes, it often generates sophisticated flow fields with convoluted solution paths (Finlay et al., 2020). This compromises the reliability of numerical solutions, resulting in highly unstable model outputs. Secondly, it is computationally expensive compared to the original FM training algorithm, for it needs repeated simulation of the ODE. These issues are major obstacles to the adoption of ODE-based models trained with MLE.

### v2

@@ 4.1 The Basic Framework
This Gaussian posterior assumption is grounded in the maximum entropy principle (Cover, 1999), and its asymptotic validity is supported by the consistency theorem (Wied and Weißbach, 2012). The following theorem specifies the MLE fine-tuning loss (proof and generalization in Appendix A).
Theorem 4.2 (MLE Training).
Under Assumption 4.1, and when $\Sigma$ is a scalar matrix, performing maximum likelihood estimation (MLE) by maximizing the expectation $\mathbb{E}_{q(x_{0},x_{1})}\left[\log p_{1}\left(\hat{\phi}_{N}(x_{0})|x_{1}\right)\right]$ is equivalent to minimizing the following loss function
@@
This forms the basic framework for fine-tuning. We only need to compute the model output $\hat{\phi}_{N}(x_{0})$, and then optimize the relevant parameters using (3). The detailed procedure is in Algorithm A.2. This enables the direct optimization of the model error, standing in contrast to previous approaches (Theorem 3.1), which could only provide a loose upper bound that contains non-optimizable components.
@@
MLE by (3) has the advantange of high-precision for directly optimizing inference results. But it suffers from several critical issues. First, it is particularly prone to overfitting. Since no additional constraints are imposed on the vector field or trajectory shapes, it often generates sophisticated flow fields with convoluted solution paths (Finlay et al., 2020), reducing the model’s reliability. Secondly, it is computationally expensive compared to the original FM training algorithm, for it needs repeated simulation of the ODE.